# MCA-Bench: A Multimodal Benchmark for Evaluating CAPTCHA Robustness Against VLM-based Attacks

### v2

@@ Appendix C More Extensive Experiments
C.1 Comparison of Model Stability and Human Performance Variability
@@
C.2 Error-Correction Capability Analysis
@@
C.3 Comparative Analysis of Micro-Decision Path Interpretability
Figure 9: Quantitative Comparison of Model Decision Paths in Multimodal CAPTCHA Scenarios. The horizontal axis includes 20 representative model architectures (e.g., 3-3G, HPR, J-S CC), with the left vertical axis indicating the normalized decision consistency index (ranging from 0 to 1.2), and the right vertical axis showing the variance fluctuation coefficient (ranging from 0 to 0.3). The red line represents the alignment of each model’s decision path with the human baseline, while the cobalt blue line reflects the degree of decision variability. Error bars denote standard deviations across three independent samples. The HPR model (Human-Pattern Recognition) demonstrates near-human performance in both consistency index (1.05 $\pm$ 0.07) and variance coefficient (0.12 $\pm$ 0.03). In contrast, the CD model (Cascade Decision) exhibits the highest decision variance (0.28 $\pm$ 0.05), revealing instability in decision logic. Notably, in the CCC (Cross-Channel Correlation) task, over 70% of models exceed the 0.2 threshold in variance coefficient.
@@ D.1 Limitations
D.2 Future Work Outlook

### v6

@@ Appendix C More Extensive Experiments
Comparison of Model Stability and Human Performance Variability
@@
Error-Correction Capability Analysis
@@
Comparative Analysis of Micro-Decision Path Interpretability
   Model    Classic-character    Distorted-word    Arithmetic-character
   Qwen2.5-VL-7B    0.320    0.985    0.695
   Human    0.920    0.840    0.985
   OCR    0.240    0.860    0.660
Table 11: Discrepancies in Recognition Performance. A Comparative Analysis of Qwen2.5-VL, Human Performance, and Traditional OCR on Classic, Distorted, and Arithmetic CAPTCHAs
Performance Comparison Between MCABench and Traditional OCR
As shown in Table 11, MCABench reveals a clear performance gap between modern multimodal models and traditional OCR systems. While OCR performs reasonably on simple distorted-word CAPTCHA, its accuracy drops sharply on classic-character and arithmetic-character types due to limited robustness against noise, rotation, and font variations. In contrast, models evaluated under MCABench—such as Qwen2.5-VL—exhibit significantly better generalization across diverse CAPTCHA structures. These results indicate that traditional OCR is increasingly inadequate for modern CAPTCHA scenarios, reinforcing the need for MLLM-based recognition approaches.
@@
Quantitative Difficulty Taxonomy Across CAPTCHA Modalities
In designing robust CAPTCHA evaluation frameworks, it is essential to systematically quantify task complexity to ensure consistent benchmarking across both human participants and automated solvers. To address this need,we introduce a unified five-level difficulty taxonomy—ranging from Easy to Very Hard—that quantitatively characterizes the complexity of four major CAPTCHA modalities: static visual recognition, point-and-click localization, interactive manipulation, and textual logic reasoning.As summarized in Table 12. By standardizing distortion strength, distractor similarity, interaction precision, reasoning depth, and other key parameters, this taxonomy establishes a fine-grained and consistent framework for benchmarking both human and model performance. The normalization enables fair cross-modal comparison and provides a practical foundation for automated difficulty adjustment and solver evaluation.
Modality Subtask Difficulty Metric Easy Medium Hard Very Hard
Static visual recognition Distorted OCR / Classic char / Color discrimination Distortion coefficient $t$ (px) $\leq 2$ 2–6 6–12 >12>12
Number of noise lines $\leq 3$ 4–10 11–25 >25>25
Contrast ratio $\geq 0.60$ 0.35–0.60 0.15–0.35 <0.15<0.15
Character size variance $\sigma_{\text{size}}$ (px) $\leq 2$ 2–6 6–12 >12>12
Point-and-click localization Grid selection / Vowel selection Target size ratio (%) $\geq 8$ 4–8 1–4 <1<1
Number of distractors $\leq 1$ 2–3 4–6 >6>6
Distractor similarity (0–1) $\leq 0.2$ 0.2–0.5 0.5–0.8 >0.8>0.8
Spatial jitter (px) $\leq 5$ 5–15 15–40 >40>40
Interactive manipulation Rotate slider / Sliding block / Jig-swap Initial deviation 5–15 / 10–30 px 15–40 / 30–80 px 40–80 / 80–200 px 80–180 / >200>200 px
Control resolution 1 / 1 px 1 / 2 px 2 / 3 px 3–5 / 5 px
Observation noise $\leq 0.5^{\circ}$ / 1 px 0.5–2 / 2–5 px 2–5 / 5–10 px $5^{\circ}$ / >10>10 px
Dynamic disturbance probability 0 0–0.1 0.1 0.3
Textual logic Q&A Arithmetic / Common sense / Multi-step reasoning Reasoning steps 1 2 3 $\geq 4$
Operand range 1–9 1–30 1–100 $\pm 1000$ / Decimal
Operation complexity Simple add/sub Add/Sub/ Mul/Div Nested operations Multi-step functions / Word problems
Distractor text ratio 0% 5–10% 10–20% $20\%$ / High ambiguity
Table 12: A Systematic Difficulty Taxonomy for CAPTCHA Modalities. Establishing Fine-Grained Metrics and Thresholds for Robust Benchmarking and Automated Solver Evaluation
Human-Centered Cross-Device Interaction Study for Behavioral Ground-Truth Acquisition
In real-world deployments, where diverse user behaviors and devices are unavoidable, capturing reliable human ground-truth is essential. As detailed in Table 13, we conducted a rigorously controlled human data acquisition study aimed at capturing fine-grained behavioral signatures across three representative interactive manipulation tasks. The participant cohort was intentionally balanced in terms of gender composition and exhibited a stable distribution of handedness, thereby minimizing potential demographic confounds in cross-condition comparisons. The resulting dataset systematically records device-dependent interaction patterns across desktop, laptop, and mobile platforms, revealing substantial variability in motor precision, input stability, and gesture controllability. Notably, the rotation block task presents a markedly elevated subjective difficulty level relative to the aligning slider and sliding block tasks, underscoring its heightened cognitive–motor coupling demands. This human-derived interaction corpus provides a critical empirical reference for calibrating task difficulty, interpreting solver behaviors, and establishing meaningful human–model performance baselines in subsequent analyses.
Task Type Participants Operating Device Handedness Difficulty
Aligning Slider 180 (M 90, F 90) Desktop (mouse drag); Laptop (touchpad slide); Mobile (single-finger slide) 14% L / 86% R 2.8
Rotation Block 180 (M 90, F 90) Desktop (mouse drag); Laptop (touchpad drag-rotate); Mobile (single-finger drag-rotate) 15% L / 85% R 3.7
Sliding Block 180 (M 90, F 90) Desktop (mouse drag); Laptop (touchpad slide); Mobile (single-finger drag) 16% L / 84% R 2.5
Table 13: Cross-Device User Interaction Characteristics and Perceived Task Difficulty. Comprehensive Analysis of Participant Demographics, Cross-Device Interaction Behaviors, and Subjective Difficulty Ratings Across Three Interactive Manipulation Task Types
LoRA Adaptation Parameter Configuration for Efficient Model Fine-Tuning
In practical model deployment, where task complexity and dataset diversity necessitate efficient adaptation, configuring low-rank adaptation modules becomes critical to achieving robust and reproducible performance. As detailed in Table 14, the LoRA parameter set collectively define the adaptation capacity, regularization behavior, and operational constraints of LoRA layers. Properly tuned, they enable efficient fine-tuning of large transformer-based models while minimizing additional computational overhead, maintaining stability, and ensuring experimental reproducibility.
Parameter Value Type Description
task_type TaskType.CAUSAL_LM Enum Specifies the LoRA task type as causal language modeling.
target_modules [”q_proj”, ”k_proj”, ”v_proj”, ”o_proj”] List[str] Indicates the Transformer attention submodules where LoRA weights are injected (Q, K, V, and output projections).
inference_mode False Bool Enables LoRA weight updates during training (non-inference mode).
r 128 Int Rank of the LoRA decomposition, determining the dimensionality of the low-rank adaptation.
lora_alpha 16 Int Scaling factor applied to adapted weights to control update magnitude.
lora_dropout 0.05 Float Dropout rate applied to LoRA layer inputs to mitigate overfitting.
bias ”none” String Specifies whether bias parameters are included in LoRA modules.
Table 14: Configuration of LoRA Adaptation Parameters. Detailed Specification of Parameter Types, Default Values, and Functional Descriptions for Low-Rank Adaptation Modules
@@ Appendix D Limitations and Future Work
Future Work Outlook
@@
Figure 9: Quantitative Comparison of Model Decision Paths in Multimodal CAPTCHA Scenarios. The horizontal axis includes 20 representative model architectures (e.g., 3-3G, HPR, J-S $C$), with the left vertical axis indicating the normalized decision consistency index (ranging from 0 to 1.2), and the right vertical axis showing the variance fluctuation coefficient (ranging from 0 to 0.3). The red line represents the alignment of each model’s decision path with the human baseline, while the cobalt blue line reflects the degree of decision variability. Error bars denote standard deviations across three independent samples. The HPR model (Human-Pattern Recognition) demonstrates near-human performance in both consistency index (1.05 $\pm$ 0.07) and variance coefficient (0.12 $\pm$ 0.03). In contrast, the CD model (Cascade Decision) exhibits the highest decision variance (0.28 $\pm$ 0.05), revealing instability in decision logic. Notably, in the CCC (Cross-Channel Correlation) task, over 70% of models exceed the 0.2 threshold in variance coefficient.